C3H8 + 5 O2 arrow 3 CO2 + 4 H2O

Calculate the mass of H2O is produced from the reaction of 6.3 g of propane.
1
Expert's answer
2020-06-22T09:17:13-0400

In the reaction

C3H8 + 5 O2 "\rightarrow" 3 CO2 + 4 H2O

1 mol of propane produces 4 mol of water:

"n(C_3H_8) = \frac{n(H_2O)}{4}"

"n(H_2O) = 4\u00b7n(C_3H_8)" .

The number of the moles of propane is the ratio of its mass "m" to its molar mass "M" (44.10 g/mol):

"n(C_3H_8) = \frac{m}{M} = \frac{6.3(g)}{44.10 (g/mol)} = 0.143 (mol)" .

Making use of the same relation to calculate the mass of water produced ("M = 18.02" g/mol):

"m = nM = 4\u00b70.143\u00b718.02 = 10.3" g.

Answer: 10.3 g of H2O is produced from the reaction of 6.3 g of propane.